Write an exponential decay function where the y-intercept is 4 and the y-values decrease by one-half as x increases by 1.
Answer by stanbon(75887)

Write an exponential decay function where the y-intercept is 4 and the y-values decrease by one-half as x increases by 1.
----
Form: y = a(b)^x
Because y-values decrease by one-half as x increases by 1
Form:: y = a(1/2)^x
---------------------------
Since y-intercept is 4,
4 = a(1/2)^0
a = 4
-----
Ans: y = 4(1/2)^x
-------------------
Cheers,
Stan H.
-------------